OpenStudy (anonymous):

I need help with this multi-step equation.7p-8=34?

OpenStudy (anonymous):

7p-8=34 7p=34+8 7p=42 p=42/7 p=6

OpenStudy (anonymous):

k thx

OpenStudy (alexandervonhumboldt2):

\(\Huge\frak\rlap{\color{blue}{Welcome~~to~~OpenStudy!!!}}{\color{orange}{\;Welcome~~to~~OpenStudy!!!}}\)